# Rapid Biomedical Research Classification: The Pandemic PACT Advanced Categorisation Engine

### v1

@@ Abstract
Omid Rohanian1,2, Mohammadmahdi Nouriborji1,7, Olena Seminog 3, Rodrigo Furst3, Thomas Mendy3, Shanthi Levanita3, Zaharat Kadri-Alab3, Nusrat Jabin3, Daniela Toale3, Georgina Humphreys5, Emilia Antonio3, Adrian Bucher6, Alice Norton3, David A. Clifton2,4 1NLPie Research, Oxford, UK 2Department of Engineering Science, University of Oxford, Oxford, UK 3Pandemic Sciences Institute, University of Oxford, Oxford, England, UK 4Oxford-Suzhou Centre for Advanced Research, Suzhou, China 5Green Templeton College, University of Oxford, Oxford, England, UK 6UK Collaborative on Development Research, London, UK 7Sharif University of Technology, Tehran, Iran omid.rohanian@eng.ox.ac.uk
@@ Acknowledgments
This work was primarily funded by Wellcome [226543] and the EDCTP2 Programme supported by the European Union. The Pandemic PACT Programme is also supported by the following grants: This research was funded by the National Institute for Health Research (NIHR) (CSA2022GloPID-R-3387) using UK Aid from the UK Government to support global health research. This work was carried out with the aid of a grant from the International Development Research Centre, Ottawa, Canada (109910-001). This work was supported by UK Research & Innovation (UKRI) under the UK Government’s Horizon Europe Guarantee under GloPID-R SEC 3 Grant Agreement no. 10061268. Whilst the funders of Pandemic PACT are engaged through the Pandemic PACT Advisory Group and have a role in the provision of funding data, they are not involved in the analysis and presentation of related findings.
This work was supported in part by the National Institute for Health Research (NIHR) Oxford Biomedical Research Centre (BRC), and in part by an InnoHK Project at the Hong Kong Centre for Cerebro-cardiovascular Health Engineering (COCHE). OR acknowledges the support of the Medical Research Council (grant number MR/W01761X/). DAC was supported by an NIHR Research Professorship, an RAEng Research Chair, COCHE, the UKRI, and the Pandemic Sciences Institute at the University of Oxford. The views expressed are those of the authors and not necessarily those of the NIHR, MRC, COCHE, UKRI, or the University of Oxford.

### v2

@@ Abstract
Omid Rohanian1,2, Mohammadmahdi Nouriborji1,7, Olena Seminog 3, Rodrigo Furst3, Thomas Mendy3, Shanthi Levanita3, Zaharat Kadri-Alabi3, Nusrat Jabin3, Daniela Toale3, Georgina Humphreys5, Emilia Antonio3, Adrian Bucher6, Alice Norton3, David A. Clifton2,4 1NLPie Research, Oxford, UK 2Department of Engineering Science, University of Oxford, Oxford, UK 3Pandemic Sciences Institute, University of Oxford, Oxford, England, UK 4Oxford-Suzhou Centre for Advanced Research, Suzhou, China 5Green Templeton College, University of Oxford, Oxford, England, UK 6UK Collaborative on Development Research, London, UK 7Sharif University of Technology, Tehran, Iran omid.rohanian@eng.ox.ac.uk
@@ Acknowledgments
This research was funded by the National Institute for Health Research (NIHR) (CSA2022GloPID-R-3387) using UK Aid from the UK Government to support global health research, as part of the EDCTP2 Programme supported by the European Union. This research was also funded by the Wellcome Trust [226543/Z/22/Z]. For the purpose of open access, the author has applied a CC-BY public copyright licence to any author-accepted manuscript version arising from this submission.
This work was also supported in part by the National Institute for Health Research (NIHR) Oxford Biomedical Research Centre (BRC), and in part by an InnoHK Project at the Hong Kong Centre for Cerebro-cardiovascular Health Engineering (COCHE). DAC was supported by an NIHR Research Professorship, an RAEng Research Chair, COCHE, the UKRI, and the Pandemic Sciences Institute at the University of Oxford. The views expressed are those of the authors and not necessarily those of the NIHR, COCHE, UKRI, NIHR, the UK Department of Health and Social Care, EDCTP, or the University of Oxford.